# Attention is All They Need: Exploring the Media Archaeology of the Computer Vision Research Paper

### v2

@@ 5. Discussion
Using a media archaeology approach, we have described the development of three aspects of the design of the contemporary computer vision research paper. First, we saw how teaser images, titles with acronyms and videos advertise the contribution of a paper, and how attention from arXiv and social media has become more important. Second, we saw how the results table was introduced for measuring the significance of a contribution, and become a ubiquitous part of the computer vision research paper. Finally, we looked at the transition from paper to PDF as a medium for disseminating research contributions and its new affordances for figure color and density. These trends have a key commonality: they make research papers easier to consume visually and more readily disposable.
@@
When we view online attention as value-producing labor, we see a clear explanation for our visual trends. Authors who write more papers which are easy to understand at a glance, easy to promote on social media, more obviously novel and more significant in the eyes of reviewers attract more attention, generating increased benefit from the labor of their peers. This benefit comes in a variety of forms including citations, interest from future students and employees and increased progress on specific research problems. Authors seeking attention can develop design innovations, like more readable tables, but the attention economy ultimately privileges large industry labs. Discussion around the CVPR social media ban echoes these attempts to prevent competition between large industry labs and graduate students. But even social media bans cannot prevent the design of the paper itself from attracting attention. In reality, the value of that attention is great enough to motivate industry labs to invest in publishing. In addition to the marketing benefits, those companies can leverage scholarly attention to signal (Feldman and March, 1981) the innovation and pedigree needed to recruit the best computer vision researchers.
We believe this pattern exemplifies a broader trend of exploiting the attention of researchers. While authors derive prestige and career progress from online attention, the researchers who read those papers do not. Instead, paying attention to the state of the discipline is essential for doing career-progressing research of their own. While staying current has always been a part of academic work, the growth of the discipline puts pressure on it, increasingly treating researchers like a crowd of interchangeable peers whose attention is needed to make new research prestigious. This process resembles proletarianization333In Marxist theory, the process by which people are assimilated into the working class, causing alienation. (Bueno, 2017, Ch. 3), and may cause the “malaise” observed by Su and Crandall (Su and Crandall, 2021). Just as these trends in computer vision were foreshadowed by trends in computer graphics, similar trends may follow in other fields. In a recent paper examining paper titles and abstracts, Ken Hyland argues that an attention economy has begun to form widely throughout academic publishing (Hyland, 2023). We encourage future study of the attention economy to focus on both the labor performed by peers, as well as the sort of visual evidence we have discussed.
@@
These problems also echo recent discussions regarding generative artificial intelligence. These technologies (some of which originate from CVPR, e.g. (Li et al., 2023)) show risk of producing online disinformation (Vykopal et al., 2023), displacing artists and writers (Jiang et al., 2023) and creating problems for future curators and historians regarding provenance and authenticity of historical images (Hazan, 2023). For example, as Jiang et al. discuss, the science fiction magazine Clarkesworld was forced to pause submissions due to excessive AI-generated submissions after ChatGPT went public in 2023 (Jiang et al., 2023). While authors submitting to computer vision conferences are not yet producing their papers using AI text generators, there are parallels regarding a deluge of content that places the burden of evaluation on the attention of reviewers. While bans on AI-generated content may slow the growth in the short-term, they do not resolve the fundamental commodification of attention, and may fail as new tools continue to blur the boundary between human and AI-generated writing.
The main implication is that ad hoc responses to symptoms of the attention economy, like banning discussion on social media or punishing low effort peer reviews will not change the pressures which are at play. Instead, computer vision is faced with a wicked problem (Buchanan, 1992), spanning peer review, hiring, promotion and tenure, in addition to academic discourse, which demands a larger-scale rethinking of the inherent tensions between facilitating scholarly communication and evaluating scholarship. Removing barriers to research communication appears to speed up the pace of research, but also changes disciplinary practices by forcing scholars into an arms race, competing for the attention of their peers.
Our inquiry also carries implications both for the design of the technologies which support publication and literature reviewing. As conferences continue to move beyond compatibility with paper, there is growing need for a LaTeX-compatible, digital-first document file format which gives authors control of the look and feel of their publication, but is machine-readable and supports accessibility tools. There is also a growing need for tools to help scholars collect and curate relevant literature. But rather than make this process as fast and frictionless as possible, we encourage work grounded in “slow design” (Odom et al., 2012) and “slow science” (Stengers, 2018) which attempts to recreate the preciousness, uniqueness and care that our participants ascribed to the papers they photocopied by hand. We also encourage, in line with the view of attention as labor, top conferences to consider compensating peer reviewers for the essential labor they perform.
@@
We also worry that these stylistic conventions may be contributing to the safety and injustice issues which currently surround machine learning (Buolamwini and Gebru, 2018; Mehrabi et al., 2021). If authors are under pressure to publish more and faster, it is easiest to do that by overstating the significance of completed work. There are a range of behaviors which contribute to this problem: from neglecting to explore the limitations of a method, to only showing favorable evaluations, to outright results fabrication. While we do not claim that authors are engaging in such behaviors, we do observe an tendency to write papers like advertisements and only minimally discuss the downsides of advertised methods. These behaviors become problematic when engineers, stakeholders or researchers in other fields who are unfamiliar with the reality of conference publishing may take the claims in papers at face value, and put minimally tested systems into production based on a general trust of computation and data.
@@ 6. Conclusion
In this paper, we have examined the commodification of researcher attention through the visual evidence in computer vision research papers. These changes lead authors to include teaser images, and the presence of teasers leads researchers to develop methods which can more easily be presented as products, ready-made for readers to download, use, and cite. Second, the results table allows researchers to demonstrate the effectiveness of their neural network methods, but then the expectation of comparison to other methods forces other kinds of papers into the same evaluation paradigm. Finally, desire for color figures led conferences to adopt digital proceedings, which has led authors to create content which cannot be viewed on paper, necessitating screen-based reading. Through Bueno, we interpret attention as a necessary form of labor in academic publishing. These trends shift that labor onto the attention of researchers and especially peer reviewers. While better tools to support paper-reading (e.g. (Qian et al., 2019; Chang et al., 2023; Head et al., 2022)) can help streamline this process and reduce the burden on individual scholars, the problem will persist unless scholarly communities address the commodification of scholarly attention. Generally, we recommend that conference organizers approach this situation as a systemic problem, and investigate measures to avoid competition between authors, slow the submission and peer review process and search for more equitable and efficient ways of distributing attention and prestige.
@@ Acknowledgements.
This work was supported in part by a NSF graduate research fellowship. The views expressed in this paper are those of the authors and do not necessarily reflect the views of the NSF or any other agencies.References

### v3

@@ 5. Discussion
Using a media archaeology approach, we have described the development of three aspects of the design of the contemporary computer vision research paper. First, we saw how teaser images, titles with acronyms, and videos advertise the contribution of a paper, and how attention from arXiv and social media has become more important. Second, we saw how the results table was introduced for measuring the significance of a contribution, and became ubiquitous. Finally, we looked at the transition from paper to PDF as a medium for disseminating research contributions and its new affordances for figure color and density. These trends have a key commonality: they make research papers easier to consume visually and more readily disposable.
@@
When we treat online attention from peers as a kind of labor—something which allows authors to generate value from the research activities of their peers—we see a clear explanation for our visual trends. Authors who write more papers which are easy to understand at a glance, easy to promote on social media, more obviously novel, and more significant in the eyes of reviewers attract more attention. Attention confers a variety of benefits including citations, interest from future mentors, students, employers, and employees, and increased progress on specific research problems. These benefits increase the value of the author’s research.
Authors seeking attention can develop design innovations, like more readable tables, but the attention economy ultimately privileges large industry labs. Discussion around the CVPR social media ban echoes these attempts to prevent competition between large industry labs and graduate students. But even social media bans cannot prevent the design of the paper itself from attracting attention. In reality, the value of that attention is great enough to motivate industry labs to invest in publishing. In addition to the marketing benefits, those companies can leverage scholarly attention to signal (Feldman and March, 1981) the innovation and pedigree needed to recruit the best computer vision researchers.
We believe this pattern exemplifies a broader trend of exploiting the attention of researchers. While authors derive prestige and career progress from online attention, the researchers who read those papers do not. Instead, paying attention to the state of the discipline is essential for doing career-progressing research of their own. While staying current has always been a part of academic work, the growth of the discipline puts pressure on readers, increasingly treating them like a crowd of interchangeable peers whose attention is needed to make new research prestigious. This process resembles proletarianization333In Marxist theory, the process by which people are assimilated into the working class, causing alienation. (Bueno, 2017, Ch. 3), and may cause the “malaise” observed by Su and Crandall (Su and Crandall, 2021). Just as these trends in computer vision were foreshadowed by trends in computer graphics, similar trends may follow in other fields. In a recent paper examining paper titles and abstracts, Ken Hyland argues that an attention economy has begun to form widely throughout academic publishing (Hyland, 2023). We encourage future study of the attention economy to focus on both the labor performed by peers, as well as the sort of visual evidence we have discussed.
@@
These problems also echo recent discussions regarding generative artificial intelligence. These technologies (some of which originate from CVPR, e.g. (Li et al., 2023)) show risk of producing online disinformation (Vykopal et al., 2023), displacing artists and writers (Jiang et al., 2023), and creating problems for future curators and historians regarding provenance and authenticity of historical images (Hazan, 2023). For example, as Jiang et al. discuss, the science fiction magazine Clarkesworld was forced to pause submissions due to excessive AI-generated submissions after ChatGPT went public in 2023 (Jiang et al., 2023). While authors submitting to computer vision conferences are not yet producing their papers using AI text generators, there are parallels regarding a deluge of content that places the burden of evaluation on the attention of reviewers. While bans on AI-generated content may slow the growth in the short-term, they do not resolve the fundamental commodification of attention, and may fail as new tools continue to blur the boundary between human and AI-generated writing.
The main implication is that ad hoc responses to symptoms of the attention economy, like banning discussion on social media or punishing low effort peer reviews, will not change the pressures which are at play. Instead, computer vision is faced with a wicked problem (Buchanan, 1992), spanning peer review, hiring, promotion, and tenure, in addition to academic discourse, which demands a larger-scale rethinking of the inherent tensions between facilitating scholarly communication and evaluating scholarship. Removing barriers to research communication appears to speed up the pace of research, but also changes disciplinary practices by forcing scholars into an arms race, competing for the attention of their peers.
Our inquiry also carries implications both for the design of the technologies which support publication and literature reviewing. As conferences continue to move beyond compatibility with paper, there is growing need for a LaTeX-compatible, digital-first document file format which gives authors control of the look and feel of their publication, but is machine-readable and supports accessibility tools. There is also a growing need for tools to help scholars collect and curate relevant literature. But rather than make this process as fast and frictionless as possible, we encourage work grounded in “slow design” (Odom et al., 2012) and “slow science” (Stengers, 2018) which attempts to recreate the preciousness, uniqueness, and care that our participants ascribed to the hardcopy papers they photocopied by hand. We also encourage, in line with the view of attention as labor, top conferences to consider compensating peer reviewers for the essential labor they perform.
@@
We also worry that these stylistic conventions may be contributing to the safety and injustice issues which currently surround machine learning (Buolamwini and Gebru, 2018; Mehrabi et al., 2021). If authors are under pressure to publish more and faster, it is easiest to do that by overstating the significance of completed work. There are a range of behaviors which contribute to this problem: from neglecting to explore the limitations of a method, to only showing favorable evaluations, to outright fabrication of results. While we do not claim that authors are engaging in such behaviors, we do observe a tendency to write papers like advertisements and only minimally discuss the downsides of advertised methods. These behaviors become problematic when engineers, stakeholders, and researchers in other fields who are unfamiliar with the reality of conference publishing may take the claims in papers at face value, and put minimally tested systems into production based on a general trust of computation and data.
@@ 6. Conclusion
In this paper, we have examined the commodification of researcher attention through the visual evidence in computer vision research papers. These changes lead authors to include teaser images, and the presence of teasers leads researchers to develop methods which can more easily be presented as products, ready-made for readers to download, use, and cite. Second, the results table allows researchers to demonstrate the effectiveness of their neural network methods, but then the expectation of comparison to other methods forces other kinds of papers into the same evaluation paradigm. Finally, desire for color figures led conferences to adopt digital proceedings, which has led authors to create content which cannot be viewed on paper, necessitating screen-based reading. Through Bueno, we interpret attention as a necessary form of labor in academic publishing. These trends shift that labor onto the attention of researchers and especially peer reviewers. While better tools to support paper-reading (e.g. (Qian et al., 2019; Chang et al., 2023; Head et al., 2022)) can help streamline this process and reduce the burden on individual scholars, the problem will persist unless scholarly communities address the commodification of scholarly attention. Generally, we recommend that conference organizers approach this situation as a systemic problem, and investigate measures to avoid competition between authors, slow the submission and peer review process, and search for more equitable and efficient ways of distributing attention and prestige.
@@ Acknowledgements.
This work was supported in part by a National Science Foundation Graduate Research Fellowship. The views expressed in this paper are those of the authors and do not necessarily reflect the views of the NSF or any other agencies.References